In Desmos, type in equations: y= 3sqaure root of x ;y=-3 square root of x


Choose all transformations from the red parent function y=3sqaure root of x to the blue transformation function y= -3sqaure root of x + 1


Options:

Down 1

Up 1

Reflective over the x axis

Right 1


1
Expert's answer
2021-03-09T00:23:51-0500

red function "y = \sqrt x"

blue function "y=-\sqrt{x} +1"



Transformations:

"y=\sqrt x \to y=-\sqrt x" (here minus "–" reflects the function over the x-axis).

"y=-\sqrt x \to y=-\sqrt x + 1" (here "+1" translates the function up by 1).


Answer:

Up 1

Reflective over the x axis